OpenStudy (anonymous):

Guys .. can u Help me with this Im not exactly sure about my solution .. Q1 : a particle moving with constant acceleration along the a axis has a velocity of 2 m/s when its position is xi=10 m. its position 5 seconds later is xf= -30 m. FIND the acceleration ..?

OpenStudy (anonymous):

i got a= 2 m/s^2

OpenStudy (anonymous):

so ...

OpenStudy (amistre64):

change in velocity, divided by time equates to acceleration the final velocity of course is change in distance, divided by change in time

OpenStudy (amistre64):

-30-10 = -40; /5 = -8 so its going 8m/s to the left it looks like; im wondering of that matters that the initial velocity was in a positive direction to begin with or are we to assume that it was relative to the direction to begin with

OpenStudy (amistre64):

if its relative to the direction to begin with, then the change from 2 to 8 is 6\[8=2+5a\]\[6/5=a\] if not, then i can see your 2